A 0.0450-π‘˜π‘” golf ball initially at rest is given a speed of 25.0 π‘š/𝑠 when a club strikes it. If the club and ball are in contact for 2.00 π‘šπ‘ , what average force acts on the ball? Is the effect of the ball’s weight during the time of contact significant? Why or why not?


1
Expert's answer
2022-02-18T08:21:52-0500

Given:

"m=0.0450\:\rm kg"

"\Delta v=25.0\:\rm m/s"

"\Delta t=0.002\:\rm s"


The average force on the ball

"F=ma=m\frac{\Delta v}{\Delta t}\\
=0.045*25.0/0.002=562.5\:\rm N"

The effect of the ball’s weight during the time of contact is significant.